# Error correction of transversal controlled-NOT gates for scalable surface code computation

### v1

@@ Abstract
Recent experimental advances have made it possible to implement logical multi-qubit transversal gates on surface codes in a multitude of platforms. A transversal controlled-NOT (tCNOT) gate on two surface codes introduces correlated errors across the code blocks and thus requires modified decoding strategies compared to established methods of decoding surface code quantum memory (SCQM) or lattice surgery operations. In this work, we examine and benchmark the performance of three different decoding strategies for the tCNOT for scalable, fault-tolerant quantum computation. In particular, we present a low-complexity decoder based on minimum-weight perfect matching (MWPM) that achieves the same threshold as the SCQM MWPM decoder. We extend our analysis with a study of tailored decoding of a transversal teleportation circuit, along with a comparison between the performance of lattice surgery and transversal operations under Pauli and erasure noise models. Our investigation works towards systematic estimation of the cost of implementing large-scale quantum algorithms based on transversal gates in the surface code.
@@ I Introduction
Quantum error correction (QEC) protects encoded logical quantum information from decoherence on the underlying physical qubits [shor_fault-tolerant_1996, calderbank_good_1996]. Recent experimental progress has led to landmark demonstrations of fault-tolerant (FT) state preparation [egan_fault-tolerant_2021, acharya_suppressing_2023, sivak_real-time_2023] , repeated error correction [schindler_experimental_2011, chen_exponential_2021, da_silva_demonstration_2024, kim_transversal_2024], and state teleportation [ryan-anderson_high-fidelity_2024] of encoded logical states. In a multitude of platforms, high-fidelity two-qubit operations are no longer strictly restricted to two-dimensional nearest-neighbor interactions [huang_comparing_2023, bluvstein_quantum_2022], opening up the possibility to implement high-rate quantum LDPC codes [panteleev_asymptotically_2022, lin_good_2022, bravyi_high-threshold_2024] and concatenated codes [yamasaki_time-efficient_2024, yoshida_concatenate_2024]. Beyond this opportunity, non-trivial connectivity can be employed for logical operations in the widely studied surface code (SC), a leading candidate for practical quantum error correction [trout_simulating_2018, bluvstein_logical_2024].
@@
In fixed-qubit architectures, the prominence of the surface code can be attributed to its 2D planar layout, nearest-neighbour connectivity, low-depth stabilization circuits, and high error tolerance threshold [fowler_surface_2012, tomita_low-distance_2014, fowler_proof_2012]. Efficient graph-based decoders, such as minimum weight-perfect matching (MWPM), perform well at correcting common circuit-level errors [wu_fusion_2023, higgott_sparse_2023]. Beyond these considerations, logical gates on surface codes are well understood and easy to implement with 2D nearest-neighbor connectivity via braiding [raussendorf_fault-tolerant_2007, fowler_surface_2012, brown_poking_2017] or lattice surgery [horsman_surface_2012, fowler_low_2018].
@@
With non-local connectivity, it is possible to implement transversal logical gates, such as the logical CNOT (tCNOT), between any pair of surface codes. As shown in Fig. 1, this requires applying physical CNOT gates between every corresponding data qubit of the control and target SC states [horsman_surface_2012]. A tCNOT creates correlated errors. For example, a bit-flip error on a data qubit in the control may be copied over to the corresponding data qubit in the target. One method to account for these error correlations, henceforth referred to as single-update decoding, is by appropriately adding syndrome history from rounds following the CNOT gate from one SC to another [dennis_topological_2002, beverland_cost_2021]. This method is suboptimal as the combined syndromes are twice as noisy as their individual components. An alternative strategy is to directly decode all the measured syndromes of the two SCs without addition, henceforth referred to as combined-hypergraph decoding. In this case, decoding based on graph algorithms cannot be used and previous works thus resort to relatively slower hypergraph decoding [cain_correlated_2024, zhou_algorithmic_2024].
Figure 1: (a) A logical transversal CNOT operation between two rotated surface codes is performed by applying physical CNOT gates between each corresponding pair of data qubits of the SC states. (b) The transversal CNOT creates correlated errors between surface codes.
@@
In this work, we benchmark the performance of the transversal CNOT gate for scalable quantum computing using three decoding methods. In particular, we study a thus-far uncharacterized strategy that we refer to as ordered decoding. In this approach, we first decode those errors in one SC state that may be copied over to the other SC state  [^1] The basic concept behind ordered decoding was mentioned in a single sentence in [beverland_cost_2021]. However, there was no concrete implementation, analytical, or numerical results provided.. We then correct for the identified errors on the second SC state before independently decoding the residual errors. Any decoder that independently locates bit- and phase-flip errors can be employed in this overall strategy; here, we use MWPM. We find that ordered decoding using MWPM is highly effective at correcting noise correlations introduced by the transversal CNOT on surface codes, outperforming the single-update and combined-hypergraph decoder in terms of thresholds. Our analysis extends beyond previously studied constant depth circuits [viszlai_architecture_2023, ryan-anderson_high-fidelity_2024, cain_correlated_2024, zhou_algorithmic_2024, wan_iterative_2024].
@@
We additionally study decoding for transversal teleportation circuits in which one of the code blocks is measured soon after a tCNOT gate. Such teleportation circuits compose a high fraction of two-qubit gate usage in quantum algorithms. We find that such teleportation operations can be inherently decoded using graph-based methods. We also provide a performance and preliminary resource overhead comparison for logical operations performed transversally versus via joint-parity measurements, i.e., lattice surgery, thus far the more widely studied method for surface code logical gate operations. This analysis is presented for Pauli-noise and a mix of erasure and Pauli noise. The latter noise model is motivated by recent studies showing that qubits with dominant erasure noise exhibit high thresholds and improved error-correction properties  [grassl_codes_1997, stace_thresholds_2009, wu_erasure_2022, kang_quantum_2023, teoh_dual-rail_2023, sahay_high-threshold_2023].
@@
Our work is structured as follows. In Section II, we provide a brief introduction to the surface code when used as a quantum memory (SCQM), followed by an analysis of methods to decode and correct errors on this system. We then move onto logical computations using surface codes. We describe the use of fault-tolerant transversal CNOTs in a general logical quantum algorithm in Section  III. In Section IV we discuss different tCNOT decoding strategies. In Section V we provide an analysis of gate teleportation, with a focus on decoding optimizations and a brief comparison with lattice surgery. Finally, in Section VI, we discuss the performance of transversal and lattice-surgery based logical gates for erasure-based noise models. We conclude in Section VII.
@@ II The surface code as an error-corrected memory
The rotated surface code [bravyi_quantum_1998, bombin_optimal_2007] is a stabilizer error correcting code [gottesman_stabilizer_1997] that uses $d^{2}$ physical qubits arranged on the vertices of a $d\times d$ square lattice to encode one logical qubit. The length of the smallest logical operator, or equivalently the minimum number of Pauli errors to cause an undetectable change in the logical state, is $d$. Here, we take $d$ to be odd. The stabilizer group $\mathcal{S}$ of this code is generated by $X$ and $Z$ type checks $S_{X}$ ($S_{Z}$) on alternating faces of this lattice, as illustrated in Fig. 2(a). Each $X$ ($Z$)-check is a product of Pauli $X$ ($Z$) operators on the qubits around the face. The $X$- and $Z$- type logical operators, $\overline{X}$ and $\overline{Z}$, consist of Pauli $X$ and $Z$ operators on qubits lying on strings connecting the boundaries of the lattice such that $\{\overline{X},\overline{Z}\}=0$, $[\overline{X},S]=0$ and $[\overline{Z},S]=0$ for all checks $S\in\mathcal{S}$. Figure 2(a) shows a distance-5 surface code with a logical operator $\overline{X}$ (marked (i)).
@@
In practice, each $X$ ($Z$) check of the rotated surface code is measured using a carefully ordered depth-4 circuit of CNOT (CZ) gates to entangle the relevant physical data qubits with an additional ancilla qubit that is then measured out [tomita_low-distance_2014], giving rise to a set of measurement outcomes referred to as the syndrome. Errors can occur within this circuit at any point. Detectable errors anticommute with a subset of checks and flip the corresponding measurement outcomes, creating defects. For example, Fig. 2(a)(ii), shows a $Z$ error on a data qubit that creates two adjacent defects. In the presence of faulty measurements, each stabilizer is generally measured $O(d)$ times [dennis_topological_2002]. This error correction protocol implements $I$ on the encoded qubit - as a result, an isolated surface code acts as a quantum memory.
@@ II.1 Efficiently decoding errors on the SCQM
Given an error syndrome $\sigma$, optimal decoding involves finding a correction that maximizes the probability of restoration to the original code state or finding the most probable logical error. For general systems, this can be a computationally hard problem [iyer_hardness_2015]. Using terminology similar to Ref. [higgott_sparse_2023], we now give an overview of simpler polynomial-time decoders used for surface code error correction.
@@
We begin by defining a decoding hypergraph $G=(V,E)$ on a surface code with errors $\mathcal{E}$. Each vertex $v^{t}_{S}\in V$ corresponds to a detector, where a detector refers to the parity between the measurement outcomes of a check at time $t-1$ and $t$. In other words, $v^{t}_{S}=S^{t-1}\oplus S^{t}$ for $S\in\mathcal{S}$, where we consider $S^{0}$ to be an initial perfect check measurement. The set of detectors with odd parity encompasses $\mathcal{D}$, the defects generated by $\mathcal{E}$. Every hyperedge $e\in E$ is a set of detectors, and is assigned a weight proportional to the logarithm of the total probability of an independent error that causes that set of detectors to take the value 1. Given $\{G,\mathcal{D}\}$, a graph-based decoder’s task is to find the most probable physical error that created $\mathcal{D}$. For general hypergraphs where hyperedges correspond to more than two detectors, this is a computationally hard problem [berlekamp_inherent_1978]. By making the simplification of finding a locally - as opposed to globally- optimal solution, it is possible to efficiently find a correction operator $\mathcal{C}$ whose syndrome matches $\mathcal{D}$; one such strategy is referred to as the hypergraph union-find (HUF) decoder [delfosse_toward_2021] .
@@
Another simplification is to reduce $G$ to a graph where $|e|\leq 2,\forall e$. For the surface code, this is possible for all single-qubit $X$ ($Z$) errors since these create independent pairs of defects (or a single defect at the boundaries) on the $Z$($X$) type stabilizer set, seen in Fig. 2(a)(ii). As a result, the decoding hypergraph $G$ can be split into two disjoint graphs, $\{G_{X},\mathcal{D}_{X}\}$ and $\{G_{Z},\mathcal{D}_{Z}\}$, that satisfy $|e|\leq 2$ $\forall e$. Fig. 2(a)(ii) shows an example of $G_{X}$ for one measurement round, and the box in Fig. 2(b) represents an example of $G_{X}$ for $d$ measurement rounds that we refer to as the spacetime decoding volume. $Y$ noise is decomposed into $G_{X},G_{Z}$ as uncorrelated $X$ and $Z$ errors occurring with the appropriate marginal probability, an approximation that performs reasonably well for locally correlated noise. A decoder can now identify the most probable physical error in polynomial time by mapping $\sigma$ to minimum-weight matching problems on $G_{X},G_{Z}$.
@@
For a given correction $\mathcal{C}$ found by a decoder, a corresponding update is applied to the surface code, ideally restoring it to the original state (as in Fig. 2(b)(iii)), but potentially causing a logical error if the correction proposed is logically inequivalent to the original error. For a MWPM-based decoder applied to a circuit-level noise model with two-qubit gate errors occurring with probability $p$, the threshold error rate is $p_{t}\approx 1\%\leavevmode\nobreak$[wang_surface_2011, gidney_stim_2021]. We find the corresponding HUF threshold to be $0.86\%$.
@@ III A scalable transversal CNOT
In a transversal multi-qubit logical operation, a physical qubit of one logical block interacts with at most one physical qubit of another logical block. This approach naturally preserves the effective code distance. Here, we focus on a logical transversal CNOT (tCNOT) between a control surface code block $C$ and a target block $T$, implemented with physical CNOT gates applied between qubit $q_{C}$ in $C$ and qubit $q_{T}$ in $T$, for every physical qubit $q$.
@@
With transversal tCNOTs, in principle it is possible to perform decoding over the entire algorithm or decode with $O(1)$ rounds of syndrome extraction using Steane- or Knill-style error correction with well-prepared ancillas [steane_active_1997, knill_fault-tolerant_2004, knill_quantum_2005, zhou_algorithmic_2024]. However, here we consider the decoding for tCNOT gates at scale. In particular we are interested in quantum algorithms with $\Theta(e^{d})$ gates that require a distance $d$ fault-tolerant gate set for successful implementation [gidney_how_2021, dalzell_quantum_2023].
@@
Consider such a classically intractable deep quantum circuit composed of many tCNOTs, where each tCNOT gate is followed by $r$ rounds of stabilizer measurements on all the code blocks. We ignore single-qubit gates for simplicity. In the following, we discuss the requirements for $r$ in the context of scalable quantum computing. We note that in this deep circuit, information from final transversal logical measurements may not be accessible until after a circuit depth of $\Theta(e^{d})$.
@@
Figure 3: A circuit composed of several tCNOTs where each tCNOT is followed by $r$ rounds of stabilizer measurement indicated by the shaded yellow box. In this case a single error (for example an $X$ error on qubit $q$, marked by a dashed red line) can induce correlated errors that grow exponentially with circuit depth (marked by dashed blue lines).
@@
It is known that for local check measurements, the surface code does not satisfy single-shot code properties [bombin_single-shot_2015, campbell_theory_2019, lin_single-shot_2023], and decoding errors at a specific location requires roughly $W=O(d)$ subsequent rounds of stabilizer measurements [dennis_topological_2002, campbell_theory_2019, delfosse_beyond_2022]. This hints at an underlying relationship between $r$ and $W$ for tCNOTs that we exemplify using the example depth-3 circuit in Fig. 3, where a single bit-flip error $X_{q}$ on a qubit $q$ in the control SC of the first tCNOT can propagate to all corresponding qubits of SCs in the circuit. In a similar fashion, $Z$ errors on the target qubits will flow to the control. Correction of this first tCNOT should address both $X$ and $Z$ errors.
@@
It is clear that when $r=W$, existing errors before each tCNOT gate can be fault-tolerantly corrected before the application of subsequent tCNOT gates. When $r\ll W$, however, one cannot decode each tCNOT independently. To see this, consider a control block of a tCNOT of the example circuit in Fig. 3, with an open $Z$-error string $E$ right before the tCNOT connecting a spatial boundary to a single defect. Given the decoding window width $r$, a minimum-weight decoder can misinterpret $E$ as a single length-$r$ measurement error string if $E$ has length $>r$. As a result, the decoder fails to correct $E$.
@@
Instead of attempting to decode individual tCNOTs it may be possible to use a correlated decoding approach when $r<W$. In this approach, the circuit is divided into an expanded depth-$m$ window that extends over all qubits, and the syndromes in this window are used to decode errors at the beginning of the window [cain_correlated_2024, zhou_algorithmic_2024]. This procedure is an extension of the overlapping- or sliding-window approach used for preserving a quantum state but applied to a quantum circuit block [skoric_parallel_2023, bombin_modular_2023, tan_scalable_2023, delfosse_spacetime_2023]. Note that some qubits may be idle in the depth-$m$ window and these can be decoded separately as conventional quantum memories. In the rest of our discussion we will neglect these idle qubits from the window.
@@
To illustrate with an example, consider the generalization of the circuit in Fig. 3 to one with depth $M$ and $2^{M}$ qubits. Suppose in order to decode $X$ and $Z$ errors in $C$ of the first tCNOT, which we label as $C_{1}$, syndromes in the depth-$m$ ($m\leq M$) window extending over $2^{m}$ (non-idle) qubits are used. In a window width of $m$, $C_{1}$ is always the control of the tCNOTs. Consequently, a $X$ error is copied from $C_{1}$ onto $2^{m}$ qubits, each of which may provide syndrome information about this propagated error. Additionally, $Z$ errors can propagate onto $C_{1}$ from $m$ logical gates. The decoding of the first tCNOT at $C_{1}$ in the presence of these $Z$ errors is dependent on decoding of the qubits that the errors originate from. The total decoding volume, which determines the complexity of decoding in the correlated decoding approach, thus scales exponentially with $m$. If the decoding volume becomes too large then the backlog problem may be amplified [terhal_quantum_2015, higgott_sparse_2023]. Thus, we need to determine how this volume scales based on a choice for $m$ and $r$ that ensure that long-lived errors are prevented.

### v2

# Error correction of transversal CNOT gates for scalable surface code computation

@@ Abstract
Recent experimental advances have made it possible to implement logical multi-qubit transversal gates on surface codes in a multitude of platforms. A transversal controlled-NOT (tCNOT) gate on two surface codes introduces correlated errors across the code blocks and thus requires modified decoding compared to established methods of decoding surface code quantum memory (SCQM) or lattice surgery operations. In this work, we examine and benchmark the performance of three different decoding strategies for the tCNOT for scalable, fault-tolerant quantum computation. In particular, we present a low-complexity decoder based on minimum-weight perfect matching (MWPM) that achieves the same threshold as the SCQM MWPM decoder. We extend our analysis with a study of tailored decoding of a transversal teleportation circuit, along with a comparison between the performance of lattice surgery and transversal operations under Pauli and erasure noise models. Our investigation builds towards systematic estimation of the cost of implementing large-scale quantum algorithms based on transversal gates in the surface code.
@@ I Introduction
Quantum error correction (QEC) protects encoded logical quantum information from decoherence on the underlying physical qubits [shor_fault-tolerant_1996, calderbank_good_1996]. Recent experimental progress has led to landmark demonstrations of fault-tolerant (FT) state preparation [egan_fault-tolerant_2021, acharya_suppressing_2023, sivak_real-time_2023, acharya2024quantumerrorcorrectionsurface], repeated error correction [schindler_experimental_2011, chen_exponential_2021, da_silva_demonstration_2024, kim_transversal_2024], and state teleportation [ryan-anderson_high-fidelity_2024] of encoded logical states. In a multitude of platforms, high-fidelity two-qubit operations are no longer strictly confined to two-dimensional nearest-neighbor interactions [huang_comparing_2023, bluvstein_quantum_2022], opening up the possibility to implement high-rate quantum LDPC codes  [panteleev_asymptotically_2022, lin_good_2022, bravyi_high-threshold_2024] and concatenated codes  [yamasaki_time-efficient_2024, yoshida_concatenate_2024]. Beyond this opportunity, non-trivial connectivity can be employed for logical operations in the widely studied surface code (SC), a leading candidate for practical quantum error correction [trout_simulating_2018, bluvstein_logical_2024].
@@
In fixed-qubit architectures, the prominence of the surface code can be attributed to its 2D planar layout, nearest-neighbour connectivity, low-depth stabilization circuits, and high error tolerance threshold [fowler_surface_2012, tomita_low-distance_2014, fowler_proof_2012]. Efficient graph-based decoders, such as minimum weight-perfect matching (MWPM), perform well at correcting common circuit-level errors [wu_fusion_2023, higgott_sparse_2023]. Further, logical gates on surface codes are well understood and easy to implement with 2D nearest-neighbor connectivity via braiding [raussendorf_fault-tolerant_2007, fowler_surface_2012, brown_poking_2017] or lattice surgery [horsman_surface_2012, fowler_low_2018].
@@
With non-local connectivity, it is possible to implement transversal logical gates, such as the logical CNOT (tCNOT), between any pair of surface codes. As shown in Fig. 1, this requires applying physical CNOT gates between every corresponding data qubit of the control and target SC states [horsman_surface_2012]. A tCNOT creates correlated errors. For example, a bit-flip error on a data qubit in the control may be copied over to the corresponding data qubit in the target. One method to account for these error correlations is by appropriately adding syndrome history from rounds following the CNOT gate from one SC to another [dennis_topological_2002, beverland_cost_2021]. Decoding using the resultant syndromes is suboptimal since the combined syndromes are twice as noisy as their individual components. An alternative decoding strategy is to directly use all the measured syndromes of the two SCs without addition. In this case, decoding based on graph algorithms cannot be used and previous works thus resort to relatively slower hypergraph decoding [cain_correlated_2024, zhou_algorithmic_2024].
@@
In this work, we benchmark the performance of the transversal CNOT gate for scalable quantum computing using three decoding methods. In addition to the approaches described above, we study a thus-far uncharacterized strategy that we refer to as ordered decoding. In this approach, we first decode those errors in one SC state that may be copied over to the other SC state  [^1] The basic concept behind ordered decoding was mentioned in a single sentence in [beverland_cost_2021]. However, there was no concrete implementation, analytical, or numerical results provided.. We then correct for the identified errors on the second SC state before independently decoding the residual errors. Any decoder that independently locates bit- and phase-flip errors can be employed in this overall strategy; here, we use MWPM. We find that ordered decoding using MWPM is highly effective at correcting noise correlations introduced by the transversal CNOT on surface codes, outperforming previous tCNOT decoders in terms of thresholds. Our analysis extends beyond previously-studied constant depth circuits [viszlai_architecture_2023, ryan-anderson_high-fidelity_2024, cain_correlated_2024, zhou_algorithmic_2024, wan_iterative_2024].
@@
We additionally study decoding for transversal teleportation circuits in which one of the code blocks is measured soon after a tCNOT gate. Such teleportation circuits comprise a high fraction of two-qubit gate usage in quantum algorithms. We find that such teleportation operations can be inherently decoded using graph-based methods. We also provide a comparison for logical operations performed transversally versus via joint-parity measurements, i.e., lattice surgery, thus far the more widely studied method for surface code logical gate operations. This analysis is presented for Pauli-noise and a mix of erasure and Pauli noise. The latter noise model is motivated by recent studies showing that qubits with dominant erasure noise exhibit high thresholds and improved error-correction properties  [grassl_codes_1997, stace_thresholds_2009, wu_erasure_2022, kang_quantum_2023, teoh_dual-rail_2023, sahay_high-threshold_2023].
Figure 1: (a) A logical transversal CNOT operation between two rotated surface codes is performed by applying physical CNOT gates between each corresponding pair of data qubits of the SC states. (b) The transversal CNOT creates correlated errors between surface codes. Each SC is shown as a qubit set on which an error on a physical qubit can propagate to the other SC through the tCNOT.
@@
Our work is structured as follows. In Section II, we provide a brief introduction to the surface code when used as a quantum memory (SCQM), followed by an analysis of methods to decode and correct errors in this system. We then move onto logical computation using surface codes. We set out definitions for tCNOT circuits and argue for individually fault-tolerant tCNOT gadgets in logical algorithms in Section  III. In Section IV we discuss different tCNOT decoding strategies. In Section V, we provide an analysis of gate teleportation, with a focus on decoding optimizations and a brief comparison with lattice surgery. Finally, in Section VI, we discuss the performance of transversal and lattice-surgery based logical gates for erasure-based noise models. We conclude in Section VII.
@@ II The surface code as an error-corrected memory
The rotated surface code [bravyi_quantum_1998, bombin_optimal_2007] is a stabilizer error correcting code [gottesman_stabilizer_1997] that uses $d^{2}$ physical qubits arranged on the vertices of a $d\times d$ square lattice to encode one logical qubit. The length of the smallest logical operator, or equivalently the minimum number of Pauli errors to cause an undetectable change in the logical state, is $d$. Here, we take $d$ to be odd. The stabilizer group $\mathcal{S}$ of this code is generated by $X$ and $Z$ type checks $S_{X}$ ($S_{Z}$) on alternating faces of this lattice, as illustrated in Fig. 2(a). Each $X$ ($Z$)-check is a product of Pauli $X$ ($Z$) operators on the qubits around the face. The $X$- and $Z$- type logical operators, $\overline{X}$ and $\overline{Z}$, consist of Pauli $X$ and $Z$ operators on qubits lying on strings connecting the boundaries of the lattice such that $\{\overline{X},\overline{Z}\}=0$, $[\overline{X},S]=0$ and $[\overline{Z},S]=0$ for all checks $S\in\mathcal{S}$. Figure 2(a) shows a $d=5$ surface code with a logical operator $\overline{X}$ (marked (i)).
@@
In practice, each $X$ ($Z$) check of the rotated surface code is measured using a depth-4 circuit of CNOT (CZ) gates to entangle the relevant physical data qubits with an additional ancilla qubit that is then measured out [tomita_low-distance_2014], giving rise to a set of measurement outcomes referred to as the syndrome. Errors can occur within this circuit at any point. Detectable errors anticommute with a subset of checks and flip the corresponding measurement outcomes, creating defects. For example, Fig. 2(a)(ii), shows a $Z$ error on a data qubit that creates two adjacent defects. In the presence of faulty measurements, each stabilizer is generally measured $O(d)$ times [dennis_topological_2002]. This error correction protocol implements the identity channel on the encoded qubit - as a result, an isolated surface code acts as a quantum memory.
@@ II.1 Efficiently decoding errors on the SCQM
Given an error syndrome $\sigma$, optimal decoding involves finding a correction that maximizes the probability of restoration to the original code state or finding the most probable logical error. For general systems, this can be a computationally hard problem [iyer_hardness_2015]. In this section, we give an overview of simpler polynomial-time decoders used for surface code error correction.
@@
We begin by defining a decoding hypergraph $G=(V,E)$ on a surface code with errors $\mathcal{E}$. Each vertex $v^{t}_{S}\in V$ corresponds to a detector, where a detector refers to the parity between the measurement outcomes of a check at time $t-1$ and $t$. In other words, $v^{t}_{S}=S^{t-1}\oplus S^{t}$ for $S\in\mathcal{S}$  [^2] We consider $S^{0}$ and $S^{d+1}$ to be initial and final perfect check measurements respectively. The defects $\mathcal{D}$ generated by $\mathcal{E}$ are detectors with odd parity under $\mathcal{E}$. Every hyperedge $e\in E$ is a set of detectors, and is assigned a weight proportional to the logarithm of the total probability of an independent error that causes that set of detectors to take the value 1. Given $\{G,\mathcal{D}\}$, a graph-based decoder’s task is to find the most probable physical error that created $\mathcal{D}$.
For general hypergraphs where hyperedges correspond to more than two detectors, this is a computationally hard problem [berlekamp_inherent_1978]. By making the simplification of finding a locally - as opposed to globally- optimal solution, it is possible to efficiently find a correction operator $\mathcal{C}$ whose syndrome matches $\mathcal{D}$; one such strategy is referred to as the hypergraph union-find (HUF) decoder [delfosse_toward_2021] .
@@
Another simplification is to reduce $G$ to a graph where $|e|\leq 2,\forall e$. For the surface code, this is possible for all single-qubit $X$ ($Z$) errors since these create independent pairs of defects (or a single defect at the boundaries) on the $Z$($X$) type stabilizer set, seen in Fig. 2(a)(ii). As a result, the decoding hypergraph $G$ can be split into two disjoint graphs, $\{G_{X},\mathcal{D}_{X}\}$ and $\{G_{Z},\mathcal{D}_{Z}\}$, that satisfy $|e|\leq 2$ $\forall e\in E$. Fig. 2(a)(iii) shows an example of $G_{X}$ for one measurement round, and the box in Fig. 2(b) represents an example of $G_{X}$ for $d$ measurement rounds that we refer to as the spacetime decoding volume. $Y$ noise is decomposed into $G_{X},G_{Z}$ as uncorrelated $X$ and $Z$ errors. A decoder can now identify the most probable physical error in polynomial time by mapping $\sigma$ to minimum-weight matching problems on $G_{X},G_{Z}$.
@@
For a given correction $\mathcal{C}$ found by a decoder, a corresponding update is applied to the surface code, ideally restoring it to the original state (as in Fig. 2(b)(iii)), but potentially causing a logical error if the correction proposed is logically inequivalent to the original error. For an MWPM-based decoder applied to a circuit-level noise model with two-qubit gate errors, the threshold error rate is $p_{t}\approx 1\%\leavevmode\nobreak$[wang_surface_2011, gidney_stim_2021]. We find the corresponding HUF threshold to be $0.89\%$.
@@ III A scalable transversal CNOT
In a transversal multi-qubit logical operation, a physical qubit of one logical block interacts with at most one physical qubit of another logical block. This approach naturally preserves the effective code distance. Here, we focus on a logical transversal CNOT (tCNOT) between a control surface code block $C$ and a target block $T$, implemented with physical CNOT gates applied between qubit $q_{C}$ in $C$ and $q_{T}$ in $T$, for every physical qubit $q$.
@@
With tCNOTs, in principle it is possible to decode over an entire algorithm using $O(1)$ rounds of syndrome extraction per tCNOT with well-prepared logical ancilla states [steane_active_1997, knill_fault-tolerant_2004, knill_quantum_2005, zhou_algorithmic_2024]. However, here we consider decoding for tCNOT gates at scale. In particular, we focus on quantum algorithms with the number of gates scaling exponentially with $d$, that require a distance $d$ fault-tolerant gate set for successful implementation [gidney_how_2021, dalzell_quantum_2023].
@@
For such a circuit, we define a tCNOT gadget to comprise of $g=O(1)$ tCNOTs in a known configuration. We take each gadget to be followed by $r$ rounds of stabilizer measurements on involved code blocks. For simplicity, we consider $g=1$, i.e. a gadget consists of a single tCNOT, and we ignore single qubit gates. In the following, we discuss the requirements for $r$ in the context of scalable quantum computing. Note that in this deep circuit, information from final transversal logical measurements is not readily accessible for use in decoding.
@@
Figure 3: (a) A ‘binary-tree’ tCNOT circuit of logical depth $M$. Each tCNOT is followed by $r$ rounds of stabilizer measurements (shaded yellow boxes). In this case a single error (for example an $X$ error on qubit $q$) can induce correlated errors that grow exponentially with circuit depth (marked by dashed red lines). (b) If $r<<W$, each tCNOT cannot be decoded independently since weight $O(r)$ data errors can be misidentified as measurement errors, creating logical failures.
@@
We exemplify our arguments using a ‘binary-tree’ circuit of logical gadget depth $M$ involving $2^{M}$ qubits. Fig. 3(a) shows an example circuit where $M=3$. In this circuit, a single bit-flip error $X_{q}$ on a qubit $q$ in the control SC of the first tCNOT, which we term $C_{1}$, can propagate to the corresponding qubits of all SCs. In a similar fashion, $Z$ errors on target SCs will flow to $C_{1}$. Correction of this first tCNOT should address both $X$ and $Z$ errors.
@@
The surface code does not satisfy single-shot code properties for local check measurements [bombin_single-shot_2015, campbell_theory_2019, lin_single-shot_2023], and decoding errors at a specific location requires roughly $W=O(d)$ subsequent rounds of stabilizer measurements [dennis_topological_2002, campbell_theory_2019, delfosse_beyond_2022]. Thus, if we set $r=W$, existing errors at each tCNOT can naturally be fault-tolerantly corrected before the application of subsequent gates. When $r\ll W$, however, one cannot decode each tCNOT independently. To see this, consider Fig. 3(b), with an open $Z$-error string $\mathcal{E}$ on $C$ right before the tCNOT connecting a spatial boundary to a single defect. Given a decoding window depth of $r$, a decoder can misinterpret $\mathcal{E}$ as a single length-$r$ measurement error string if $|\mathcal{E}|>r$. As a result, the decoder fails to correct $\mathcal{E}$.
@@
When $r<W$, instead of attempting to decode individual tCNOTs, it may be possible to use a correlated decoding approach. Here, we use an expanded depth-$m$ ($m\leq M$) window of the circuit extending over $2^{m}$ qubits. Syndromes in the entire window are used to decode errors at the beginning of the window [cain_correlated_2024, zhou_algorithmic_2024]. This procedure is an extension of the overlapping- or sliding-window approach used for preserving a quantum state but applied to a quantum circuit block [skoric_parallel_2023, bombin_modular_2023, tan_scalable_2023, delfosse_spacetime_2023]. Note that in a general circuit, some qubits may be idle in the depth-$m$ window and these can be decoded separately as conventional quantum memories. In the rest of our discussion we will neglect these idle qubits.
$r$ $m$ tCNOTs decoded $C_{1}$ is Decoding
independently FT volume
$O(1)$ $O(1)$ N N $O(1)$
$O(1)$ $O(d)$ N Y $O(e^{d})$
$O(d)$ $O(1)$ Y Y $O(d)$
Table 1: Constraints limiting tCNOT gadget separation $r$ and decoding window depth $m$ while decoding an exponentially growing tCNOT circuit. We assume $O(1)\ll d$.
@@
Let us examine correction of $X$ and $Z$ errors on $C_{1}$ in this window. Since $C_{1}$ is always the control of the tCNOTs, any error $X_{q}$ is copied from $C_{1}$ onto $2^{m}$ qubits, each of which may provide syndrome information about this propagated error. Conversely, $Z$ errors propagate onto $C_{1}$ from $m$ logical gates. Decoding of $Z$ errors on $C_{1}$ at the first tCNOT is dependent on decoding of the $2^{m}$ SCs that these errors can originate from. The total decoding volume for $C_{1}$, which determines the complexity of decoding in the correlated decoding approach, thus scales exponentially with $m$. If the decoding volume becomes too large, the backlog problem may be amplified [terhal_quantum_2015, higgott_sparse_2023]. Thus, we need to determine how this volume scales based on an $(r,m)$ choice that ensures that long-lived errors are prevented.